OpenStudy (mic123):

determine whether y varies directly with x. if so find the constant of variation k and write the equation. x 4,7,10,13 y 6.4, 11.2, 16, 20.8 Please help

OpenStudy (welshfella):

y = kx where k is the constant of variation k = y/x now work out values of k from the given data k = 6/4 = 1.5 for the first pair of data check out the other values to see if they also equal 1.5

OpenStudy (welshfella):

if so we have direct variation.

OpenStudy (welshfella):

Oh sorry i misread the data the first pair is 6.4 / 4 not 6/4

OpenStudy (mic123):

what do I do to see if the other numbers = 1.5?

OpenStudy (welshfella):

which makes k = 1.6

OpenStudy (mic123):

or 1.6

OpenStudy (welshfella):

next one to check is 11.2 / 7

OpenStudy (welshfella):

OK?

OpenStudy (mic123):

I din't understand can you please explain a little more

OpenStudy (mic123):

don't*

OpenStudy (welshfella):

for it to be direct variation the value of y/x must be the same for each pair of data ( x value and corresponding y value)

OpenStudy (welshfella):

eg when x = 7 , y = 11.2 y/x = 11.2 / 7 = 1.6 that is the same as for the first pair (4,6.4) test the other 2 values of x and y in same way

OpenStudy (welshfella):

10,16 and 13,20.8 need to be checked

OpenStudy (welshfella):

is that clearer?

OpenStudy (mic123):

a lil

OpenStudy (mic123):

I'm horrable with math

OpenStudy (welshfella):

well work out 16/10 and 20.8/ 13 and see if they come to 1.6

OpenStudy (welshfella):

if they do then you've proved direct variation

OpenStudy (welshfella):

use your calculator of course

OpenStudy (mic123):

16/10 comes out as 8/5 but if I do 1.6/10 it goes to 0.16 also 20.8/13 is 1.6

OpenStudy (welshfella):

16 / 10 = 1.6

OpenStudy (mic123):

oh ok

OpenStudy (mic123):

so is that my answer?

OpenStudy (welshfella):

so now ee its direct variation and k = 1.6 so the equation y = kx becomes y = 1.6x the required equation is y = 1.6x

OpenStudy (mic123):

cool thanks so much

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

gotta go

OpenStudy (mic123):

ok